# Smallest Suffixient Sets as a Repetitiveness Measure

### v1

@@ 5.3 Uncomparability of $\chi$ with copy-paste measures
Finally, we show that $\chi$ is incomparable with most copy-paste measures. This follow from $\chi$ being $\Theta(n)$ on de Bruijn sequences and $O(1)$ on Fibonacci strings. Because $g=O(n/\log n)$ on de Bruijn sequences [27] and by Lemma 5, we have:
@@
From Corollary 4 and Lemma 12, and that $z$, $z_{no}$, $z_{end}$, $v$, $g_{rl}$ and $c$ are always $O(g)$, it follows that there are string families where $\mu=o(\chi)$, for any $\mu\in\{z,z_{no},z_{e},z_{end},v,g,g_{rl},c\}$. On the other hand, from Lemma 11 and Corollary 3, and that $c=\Omega(\log n)$ on Fibonacci words [27, Thm. 32] and $c=O(\mu)$ for any $\mu\in\{z,z_{no},z_{e},z_{end},g_{rl},g\}$ [27, Thm. 30], it follows that there are string families where $\chi=o(\mu)$, for any $\mu\in\{z,z_{no},z_{e},z_{end},v,g,g_{rl},c\}$.∎
@@ 6 Conclusions and Open Questions
We have contributed to the understanding of $\chi$ as a new measure of repetitiveness, better finding its place among more studied ones. Figure 1 shows the (now) known relations around $\chi$ (cf. [26]).
@@
Regarding edit operations, it seems that that $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)$ is $O(1)$ for all the string operations we considered. Showing a multiplicative constant for insertion would imply the existence of a constant for rotation and vice versa. It is also open whether $r=O(\chi\log\chi)$. If this were true —and provided that $\chi$ has $O(1)$ multiplicative sensitivity to string operations— it would imply that $r$ has $O(\log n)$ multiplicative sensitivity to these operations, making the already known lower bounds on multiplicative sensitivity [1, 14, 15] tight. If the conjecture were false, then $\chi$ could be considerably smaller than $r$ in some string families.
@@ Acknowledgements
We thank Davide Cenzato for his code to compute smallest suffixient sets, which was helpful to analyse the behavior of $\chi$.

### v2

@@ 5.3 Uncomparability of $\chi$ with Copy-Paste Measures
Finally, we show that $\chi$ is incomparable with most copy-paste measures. This follows from $\chi$ being $\Theta(n)$ on de Bruijn sequences and $O(1)$ on Fibonacci strings. Because $g=O(n/\log n)$ on de Bruijn sequences [28] and by Lemma 5, we have:
@@
From Corollary 5 and Lemma 12, and that $z$, $z_{no}$, $z_{end}$, $v$, $g_{rl}$ and $c$ are always $O(g)$, it follows that there are string families where $\mu=o(\chi)$, for any $\mu\in\{z,z_{no},z_{e},z_{end},v,g,g_{rl},c\}$. On the other hand, from Lemma 11 and Corollary 4, and that $c=\Omega(\log n)$ on Fibonacci words [28, Thm. 32] and $c=O(\mu)$ for any $\mu\in\{z,z_{no},z_{e},z_{end},g_{rl},g\}$ [28, Thm. 30], it follows that there are string families where $\chi=o(\mu)$, for any $\mu\in\{z,z_{no},z_{e},z_{end},v,g,g_{rl},c\}$.∎
6 Online Computation of Smallest Suffixient Sets
As an application of Lemma 2, we show that Ukkonen’s linear-time online construction of suffix trees [33] can be easily modified to compute smallest suffixient sets within the same space and time complexity.
The suffix tree of a text $T$ is a tree of size $O(|T|)$ where edges are labeled by nonempty substrings of $T$ (the label $T[i\mathinner{.\,.}j]$ is indicated by the pair $(i,j)$). Every node has at least two children, and the labels of edges to any two children must differ in their first symbol. If node $x$ has a child node $y$ by an edge labeled $(i,j)$, we say that $y$ is a child by label $T[i]$. Each node $x$ is said to be the locus of the string obtained by concatenating the labels of the edges that lead from the root to $x$. The string depth of a node is the length of its locus. The tree leaves are the loci of suffixes $T[i\mathinner{.\,.}]$, whereas internal nodes are loci of strings that occur at least twice in $T$. Edges to leaves have labels of the form $(i,\infty)$, which means the second component is always the last position processed of $T$. Suffix tree nodes also have so-called suffix links, which lead from a node that is the loci of a string $cu$, for $c\in\Sigma$, to the loci of string $u$ (which always exists: if $cu$ occurs more than once, so does $u$). Finally, we extend the concept of suffix tree nodes to implicit nodes, which are virtual nodes with one child, assumed to exist along edges: if $y$ is the child of $x$ by an edge labeled $(i,j)$ and $x$ is the locus of string $u$, then $(x,(i,p))$, for $i\leq p<j$, is an implicit nodes whose locus is $u\cdot T[i\mathinner{.\,.}p]$. Knowing their represented string, suffix links are also defined from implicit nodes (those can lead to explicit or to implicit nodes).
Ukkonen’s algorithm builds the suffix tree of $T$ such that, after having processed any prefix $w$ of $T$, it has built the suffix tree of $w$. To prevent special cases, it assumes that the root is in fact a child of a special node $\bot$, with edges to the root labeled $c$ for every $c\in\Sigma$; the suffix link of the root is $\bot$. We do not aim at a full explanation of the algorithm, but just highlight some of its relevant properties. Let the prefix $w$ of $T$ be followed by $c\in\Sigma$. The algorithm maintains pointers to two (possibly implicit) nodes of the suffix tree of $w$:
$s$:
called the active point of $w$, is the locus of the longest suffix $u$ of $w$ that occurs at least twice in $w$;
$s^{\prime}$:
called the end point of $w$, is the locus of the longest suffix $v$ of $w$ such that $vc$ occurs in $w$.
Note that this implies that the locus of $vc$ is the active point of $wc$ (i.e., the new $s$ after processing $c$), and that, because $v$ must be a suffix of $u$, there is a chain of consecutive suffix links from $s$ to $s^{\prime}$. The algorithm updates the suffix tree nodes in that chain, from $s$ to (but not including) $s^{\prime}$, by adding a new leaf child labeled $c$ to those nodes. Although updating the suffix tree of $w$ to obtain that of $wc$ may take non-constant time, the time does amortize to constant [33].
We will enhance the suffix tree nodes by adding a mark to the nodes that are loci of the supermaximal right-extensions of a smallest suffixient set of $w$. The length of the extension is the string depth of the marked node, and any suffix descending from the node serves as a starting position of such extension. This is the way in which we maintain a smallest suffixient set for the current prefix $w$ of $T$. For example, we can easily maintain the marked nodes in a list in order to collect the set for any prefix in optimal time. To compute $\chi(T)$, we can run the algorithm on $T\text{\tt\textdollar}$ and return the length of the list.
Note that the loci of right-extensions are either explicit nodes, or implicit nodes of the form $(x,(i,i))$, because they extend by one symbol a string that appears more than once. We can then associate the mark to the corresponding explicit child $y$ of $x$, so as to consult and update it in constant time.
The first letter $c$ of $T$ is processed by adding a child leaf of the root by label $c$ and setting $s$ to the root. From there on, considering the cases of Lemma 2, we proceed as follows:
$s$ has a child by label $c$:
This means that $s=s^{\prime}$ and Ukkonen’s algorithm just descends by $c$ to find the new $s$. We do not need any further action because we are in case 1 of the lemma.
$s$ has two or more children, none by label $c$:
This is case 2 of the lemma. Ukkonen’s algorithm will create a new leaf child by label $c$ of $s$, and of all the nodes in the suffix-link chain until it finds $s^{\prime}$ (i.e., the first node in the chain having a child by label $c$). We then (1) mark the (just created) child of $s$ by label $c$, and (2) unmark, if it is marked, the (already existing) child of $s^{\prime}$ by label $c$. This is because $uc$ is a new supermaximal right extension of the right-maximal string $u$ and, in case (2), $vc$ ceases to be a supermaximal right extension because it is a suffix of the new one we are adding, $uc$.
$s$ has just one child by label $a\neq c$:
This is case 3 of the lemma. Ukkonen’s algorithm proceeds exactly as in the previous case, finding the first $s^{\prime}$ with a child by label $c$ in the suffix-link chain. We (1) mark the two children of $s$ (by labels $a$ and $c$), (2) unmark, if it is marked, the child of $s^{\prime}$ by label $c$, and (3) unmark, if it is marked, the child of $s^{\prime\prime}$ by label $a$, where $s^{\prime\prime}$ is the first node in the suffix-link chain that has another child labeled $b\neq a$ (it might be that $b=c$ and thus $s^{\prime\prime}=s^{\prime}$). This is because $ua$ and $uc$ are new supermaximal right-extensions of $u$, and in case (2), $vc$ is not anymore supermaximal, as before. In case (3), similarly, if $s^{\prime\prime}$ is the locus of $z$, then $za$ is not anymore supermaximal. Note that the child by label $c$ of $s^{\prime\prime}$ is the only locus of some $za$ suffix of $ua$ that could possibly be marked.
Ukkonen’s algorithm is claimed to be $O(n)$ time, but this assumes that the alphabet is constant. If it is not, we may need more time to find the child by label $c$ among its children, or determine it does not exist. If the alphabet is an integer range and of size polynomial in $n$, we can still retain $O(n)$ time in the transdichotomous RAM model of computation with computer word size $w=\Omega(\log n)$ using fusion trees, as the children operations can then be handled in time $O(\log_{w}|\Sigma|)=O(1)$ [31]. Otherwise, an extra factor of $O(\log|\Sigma|)$ appears.
Theorem 6.1
There exists an algorithm to compute smallest suffixient sets that processes a text $T$ left to right such that, for every $n$, after having processed the prefix $T[1\mathinner{.\,.}n]$ it has determined a smallest suffixient set of $T[1\mathinner{.\,.}n]$ using $O(n)$ space and $O(n)$ worst-case time. This can be used to compute $\chi(T)$ within the same complexities, using $n=|T|$.
@@ 7 Conclusions and Open Questions
We have contributed to the understanding of $\chi$ as a new measure of repetitiveness, better finding its place among more studied ones. Figure 1 shows the (now) known relations around $\chi$ (cf. [27]). As a direct consequence of our findings, we derive a new simple online algorithm to compute smallest suffixient sets, and thus $\chi$, as a modification of Ukkonen’s suffix tree construction [33].
@@
Regarding edit operations, it seems that $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)$ is $O(1)$ for all the string operations we considered. Showing a multiplicative constant for insertion would imply the existence of a constant for rotation and vice versa. It is also open whether $r=O(\chi\log\chi)$. If this were true —and provided that $\chi$ has $O(1)$ multiplicative sensitivity to string operations— it would imply that $r$ has $O(\log n)$ multiplicative sensitivity to these operations, making the already known lower bounds on multiplicative sensitivity [1, 14, 15] tight. If the conjecture were false, then $\chi$ could be considerably smaller than $r$ in some string families.
@@ Acknowledgements
We thank Davide Cenzato, Nicola Prezza, and Francisco Olivares for their code to compute smallest suffixient sets https://github.com/regindex/suffixient [5], which was helpful to propose and discard hypotheses on the behavior of $\chi$, and for useful discussions on suffixient sets.
Disclosure of Interests
The authors have no competing interests to declare that are relevant to the content of this article.
Funding
G.N. and C.U. were partially funded by Basal Funds FB0001 and AFB240001, ANID, Chile; and FONDECYT Project 1-230755, ANID, Chile.
G.R. was partially funded by the MUR PRIN Project “PINC, Pangenome INformatiCs: from Theory to Applications” (Grant No. 2022YRB97K), funded by Next Generation EU PNRR M4 C2, Inv. 1.1 and by the INdAM - GNCS Project CUP_\_E53C24001950001.
C.U. was partially funded by ANID-Subdirección de Capital Humano/Doctorado Nacional/2021-21210580, ANID, Chile; and NIC Chile Doctoral Scholarship, NIC, Chile.